OpenStudy (anonymous):

What is the missing term, x, in the pattern below? 1, 16, x, 256, 625, …

OpenStudy (anonymous):

Hint: Write each term as something raised to the 4th power

OpenStudy (anonymous):

81

OpenStudy (anonymous):

1^4 = ? 2^4 = ? 3^4 = ? 4^4 = ? ...

OpenStudy (anonymous):

yes